What is the angle measurement(distance) of a rectangle that is 26ft by 38ft. (Corner to corner.
Thank you
Dan

Answer by nerdybill(7384)

26^2 + 38^2 = h^2
676 + 1444 = h^2
2120 = h^2
46.043 = h
or, hypotenuse is
46 feet